# EdgeGaussians - 3D Edge Mapping via Gaussian Splatting

### v1

@@ 3.3 Edge Fitting
Given a 3D edge point cloud, there are mainly two edge fitting strategies: the fit-then-cluster approach used in NEF [91] that operates in a multi-RANSAC [20] fashion. The clusters are simply the inlier supports of each edge. The second strategy is the cluster-then-fit approach, also used in EMAP [44]: it first clusters the points belonging to the same edges and then fits an edge to each cluster. We adopt the latter strategy as the qualitative results of NEF [91] show that the first strategy sometimes ignores the local geometry of the 3D edge points. For example, curves get fitted by lines, and neighboring points belonging to the same edge are inliers to different edges.
Clustering points into edges.  A common strategy to cluster points is to adopt a graph-traversal approach where points form the vertices of the graph and an edge exists between two vertices only if they are spatially close [55, 22]. Clustering the points into their supporting edges then amounts to solving a graph traversal problem. This simple solution is usually made robust by integrating a smoothness constraint between neighboring points [57]: two vertices are connected not only if they are spatially close but if their direction is also similar. This is what motivates the two constraints on the Gaussian’s orientation previously described in Eq. 4 and  Eq. 5. We adopt this graph traversal approach that is also used in EMAP [44] and has proven to be effective.
Given the trained Gaussians, we define the graph as follows. The vertices of the graph represent oriented edge points: the vertex’s position is the Gaussian’s center and the vertex’s direction is the Gaussian’s principal direction. Vertices are neighbors if they are spatial neighbors and if they have similar orientations. During traversal, we add a third orientation criterion for robustness: a new vertex is added to the current path / cluster only if its orientation is consistent with the path’s global direction, i.e., the edge direction. This is measured by comparing the direction of the path and the direction between the new vertex and the one last added to the path. These two directions should be close. This favors sharper clustering around corners where the learned direction can get smooth. For all orientation tests, we use a single orientation threshold $\theta$. The graph traversal results in clusters of points belonging to the same edge. We then fit edges to the resulting clusters in the form of line segments and cubic Bezier curves.
@@
Parametric Edge Fitting.  The previous clustering generates groups of points that belong to the same edges on top of which we estimate a parametric representation of the edges. These edges can be either 3D lines or 3D curves, each having a different parametric representation so one must select which parametric model to fit on each cluster. To select between the two, we fit both a line and a curve on a given cluster and use the residuals’ error to select the best model.
Let $e_{\text{c}}$ and $e_{\textit{l}}$ be the average residual error of the curve and line models respectively. We select the curve model if the error $e_{\text{c}}$ is smaller than a fraction $\delta$ of the line residual error, i.e., $e_{\text{c}}\leq\delta e_{\textit{l}}$. The parameter $\delta$ can be tuned to control the fraction of lines and curves in the output. As shown in the qualitative results (Fig. 4), this strategy is more effective than the ones adopted in NEF [91] and EMAP [44]. NEF [91] first approximates all edges with 2-control-points Bezier curves, i.e., 3D lines, and uses the estimated control point to further optimize the models into 3D curves when needed. This approach produces several degenerate curve configurations, especially close to the corners. As for EMAP [44], it first fits 3D lines, filters out their inlier points, then fits 3D curves on top of the remaining points. This has the disadvantage of making the line fitting agnostic to the presence of curves so a large fraction of curves get fitted by lines instead.
@@ 4.1 Experimental Setup
Datasets.  We follow the quantitative evaluation defined in EMAP [44] that runs on two datasets: ABC-NEF [91, 41] and DTU [35]. The evaluation runs on the data released by EMAP [44], which includes 2D edge maps [73, 62] used for training, and with their released evaluation code. We also report qualitative results on the Replica dataset [72] in the supplementary material.
@@
The ABC-NEF [91] dataset is a subset of the ABC dataset [41] curated specifically for the task of 3D edge mapping. It holds 115 CAD models, ground-truth parametric edges, and 50 views rendered from around the object. We use the subset of ABC-NEF selected by the authors of EMAP [44] that keeps 82 out of the 115 CAD models obtained after filtering out models with cylinders and spheres - ones that are inconsistent with edge observations. The DTU dataset [35] comprises multi-view images of everyday objects captured from fixed frontal viewpoints. We evaluate on the 6 DTU scenes for which the EMAP [44]’s authors provide pseudo-ground-truth 3D edge points. The pseudo-ground-truth is obtained by projecting 3D points measured by a structured-light scanner [1] onto the images and by filtering out the points that do not project to 2D edges [10]. Although we perform at par or better than the baselines on DTU, we observe that the pseudo-ground-truth introduces biases in the evaluation and we illustrate them in the qualitative results. Finally, as in [44], we report qualitative results on three scenes of the Replica dataset [72] that comprises high-quality reconstructions. We do not run quantitative evaluation on Replica for it does not provide any ground-truth edges.
Baselines.  We evaluate the line-SfM-based method LIMAP [46] that sets the state-of-the-art in geometry-based 3D line reconstruction and the 3D wireframe learning method NEAT [87]. We also evaluate the state-of-the-art 3D line / curve mapping methods NEF [91] and EMAP [44]. Following [44], NEAT [87] is not evaluated on ABC-NEF [91]: the authors observe that it often fails to train on the textureless renderings of the CAD models. The qualitative results we report are generated with the code and the weights released by the authors of the various baselines. For EMAP [44], at the time of writing, running the released code with the recommended configurations does not yield the exact same results as the ones shown in the paper or the website. We therefore report the results run on checkpoint versions graciously provided by the authors. The quantitative results for the baselines are reported from [44].
Metrics.  The ground-truth edges for the ABC-NEF [91] dataset are provided so parametric edges can be evaluated quantitatively. The evaluation defined in [44] samples points along the predicted parametric edges and compares those points against points sampled at the same resolution on the ground-truth edges to compute the metrics. The accuracy defines the mean distance from the predicted points to the closest ground-truth points, and the completeness defines the mean distance between the ground-truth points and their nearest predicted point. For these two metrics, the lower the better. The precision at a distance threshold $\tau$ (P($\tau$)) measures the percentage of predicted points with at least one ground-truth point within distance $\tau$. Symmetrically, the recall (R($\tau$)) measures the percentage of ground-truth points for which a predicted point is within the distance threshold $\tau$. For precision and recall, the higher the better. We report the metrics for $\tau$ in 5, 10, and 20 millimeters (mm). Note that the used error thresholds are quite strict, as ABC-NEF [91] objects are scaled so that the largest edge of the bounding box measures 1 meter (m). We also compare the running times of the different neural implicit representation methods to show the advantage of the explicit representation adopted in this paper. We do not report the Edge Direction Consistency metric (Norm) [44] as, at the time of the writing, the theoretical definition and the implementation of that metric are not available.
Although these metrics provide a meaningful evaluation of the reconstruction quality, they do not account for all the performance aspects of the 3D edge reconstruction. For example, duplicate edge reconstructions that lie close to a ground-truth edge are not penalized. This is especially important on the DTU dataset which pseudo-ground-truth 3D edge points do not lie exactly on the 3D edges but along the 3D edges, as illustrated in Fig. 5.
Figure 5: Qualitative results and pseudo-ground-truth on DTU [35]. We visualize our method (red) and the pseudo-ground-truth edge points (green) generated in [44] by projecting ground-truth 3D points onto an image and filtering the ones that fall out of 2D edges [10]. Note that our method gets penalized in precision when it predicts geometrically correct 3D edges that are missing from the pseudo-ground-truth, e.g., the top of the smurf and the howl. Also, the pseudo-ground-truth contains thick patches of points that make it difficult for well-predicted edges to get a high recall. Still, our method faithfully reconstructs the 3D parametric edges of the objects.
[…]
4.2 Results
Figure 6: Qualitative results on DTU [35]. Comparison of the proposed 3D reconstruction method against baselines. NEAT [87] produces partially complete reconstructions with lines only. The EMAP [44] edges are more complete but there are duplicate edge predictions for a single target edge in 3D, while our reconstruction is much cleaner with mostly a single predicted edge boundary (see zoom-ins).
@@ 4.2 Results
The accuracy and the completeness of our method are on par with EMAP [44] and significantly better than NEF [91]. However, the accuracy of all learning-based methods falls behind the accuracy of the geometry-based LIMAP [46]: it predicts points closer to the ground-truth ones although its completeness has room for improvement. This is indicated by the relatively lower completeness and recall metrics. Note that the ABC-NEF [91] CAD models are scaled to fit the unit cube so objects span one meter, which makes the millimeter (mm) error thresholds used in the evaluation extremely harsh. LIMAP [46] demonstrates impressive results followed by EMAP [44] that performs better than the proposed method under the 5mm error threshold although this trend is inverted under the 10mm and 20mm error thresholds. We note that the performance of the proposed method when supervised with the 2D edge maps generated by the PidiNet [73] detector is slightly worse than with DexiNed [62], although it remains very competitive with all baselines.
We explain the lower performance of our method under the 5mm error threshold by the bias that exists in the 2D edge maps used as a supervisory signal during training. This bias is introduced by the off-the-shelf detectors [73, 62] that generate 2D edges typically thicker than the projection of the ground-truth 3D edges (See example in Fig. 5). The learned Gaussians remain valid as long as they project onto the 2D edge maps, which is the case for 3D Gaussians lying close to but not on the 3D edges. In fact, the thicker the 2D edges, the farther away from the 3D edges can the valid Gaussians be. We verify this hypothesis visually and observe that even edge reconstructions with poor metrics exhibit 3D edge points that are close to the ground-truth edge. Independently of the reported results, we approximately quantify this bias by fitting the predicted 3D edge points to the ground-truth one and measure that the 2D edge map bias induces an error in the 3D Gaussian position of $\sim$0.5of the scene size, which is $\sim$5mm for the 1 meter-wide objects in the ABC-NEF [91].
@@
Evaluation on DTU [35] The quantitative results are reported in Tab. 2-left and the qualitative results in Fig. 5 and Fig. 6. On each of the 6 DTU scans, our method either outperforms or is at par with the baselines. However, we observe biases in the evaluation introduced by the pseudo-ground-truth generation process. As mentioned in Sec. 4.1, the pseudo-ground-truth is obtained from a 3D point cloud by filtering out the points that do not project to edges on the 2D edge maps [10]. However, the edge maps exhibit edges [73, 62] thicker than the projection of the actual ground-truth 3D edges (Fig. 5). So the 3D points that get labeled as pseudo-ground-truth edge points are not only the points lying on the ground-truth edges but also the points within an $\epsilon$-bound on the edge, and the range of the bound is a function of the 2D edge map thickness and the distance between the object and the camera. We refer to this bias as the thickness bias. Another bias originates from the occlusions in the 2D edge maps that cause some 3D points to be incorrectly labeled as non-edge points. Fig. 5 illustrates these biases and compares the pseudo-ground-truth points (green) to the points sampled on the parametric edges predicted by our method (red). It can be clearly seen that our parametric edges are more complete than the pseudo-ground-truth points. Several predicted edges, for example in the hat and legs of the smurf soft toy, are not covered by the ground-truth, which illustrates the occlusion bias. This bias hinders the precision of our method even though the predictions are correct. The thickness bias is illustrated by the relatively thicker pseudo-ground-truth curves in green compared to the thinner edge predictions in red. This bias not only penalizes the recall of our method but also incorrectly favors methods that reconstruct duplicate edges along the ground-truth edges, which is undesirable as it makes the reconstruction noisier. Fig. 6 illustrates the edges predicted by our method and the baselines and show examples where EMAP [91] predicts duplicate edges whereas our method generates single and precise edges. As for the runtime, our method is as efficient as on the ABC-NEF [91] dataset and runs 22 times faster than NEF [44] and more than a hundred times faster than EMAP [44] (Tab. 2-right).
@@ 5 Conclusion
In this work, we propose a method for 3D edge reconstruction from images that explicitly learn the 3D edge points on top of which the 3D edges are fitted. The explicit representation is simple and casts 3D edge points as 3D Gaussians and the edge direction as the principal axis of the Gaussians. Such a representation allows for efficient rendering-based training supervised with off-the-shelf 2D edge maps. Results show the efficiency of the proposed explicit representation that enables 3D edge reconstruction performances on par with previous implicit representations yet runs an order of magnitude faster. While the off-the-shelf 2D edge maps make for a relevant supervisory signal, we observe that they can introduce bias in the training or the evaluation, which calls for investigating better supervision in the future.
Acknowledgements
This work was supported by the Czech Science Foundation (GACR) EXPRO (grant no. 23-07973X), the Chalmers AI Research Center (CHAIR), WASP and SSF. The compute and storage were partially supported by NAISS projects numbered NAISS-2024/22-637 and NAISS-2024/22-237 respectively.
@@ Acknowledgements
Appendix A provides implementation details about the training of the edge-specialized Gaussian Splatting in addition to those provided in the main paper. Appendix B shows qualitative results over the test split of the Replica [72] dataset defined by the authors of EMAP [44]. Fig. 10 illustrates the limitations and failure cases of our method.
Appendix A Additional Implementation details
Gaussian Parameters Initialization.  At the beginning of the training, the Gaussians are initialized as follows. Gaussian Position: The CAD models of the ABC-NEF [91] dataset comprises texture-less objects on which SfM [70] generates extremely sparse or no point reconstruction at all. We therefore initialize the Gaussian means with 10000 randomly sampled points inside a unit cube. For scenes from the DTU [35] and Replica [72] datasets, we use the SfM [70] points as initialization. We note that the method also works with similar accuracy with the random point initialization. On the Replica [72] scene, the set of points is larger as the scene is larger and contains more elements. Scenes across all datasets are scaled so that the scene fits in a box which largest side is 1 meter (m). Gaussian Scale: We use a constant initial value of 0.004 for all datasets. Gaussian Opacity: We use a constant initial value of 0.08 for all datasets. Gaussian Orientation: Random unit quaternions are used as initial values for all Gaussians.
Training Strategy.  We train the model for 500 epochs. For the first 30 epochs, we only train the position parameters so that the scale and the orientation of the Gaussian do not compensate for its incorrect position during rendering. Thus the training constrains the Gaussian’s position, i.e.mean, to lie on 3D edges. We cull the Gaussians based on opacity and duplicate the ones with high positional gradients at regular intervals similar to the original work [38]. The learning rates of different parameters are as follows. Position: starting with $1e^{-3}$, scaled with a factor of 0.75 every 10 epochs, 5 times. Scale: $2e^{-4}$ constant. Opacity: $3e^{-2}$ constant. Orientation: $1e^{-3}$ constant.
Appendix B Qualitative Results over Replica

### v2

@@ 3.3 Edge Fitting
Given a 3D edge point cloud, there are mainly two edge fitting strategies: the fit-then-cluster approach that operates in a multi-RANSAC [17] fashion. The clusters are simply the inlier supports of each edge. The second one is the cluster-then-fit approach in which points are first clustered into individual edges. The first one is used in NEF [83] while the second one is used in EMAP [41]. The fit-then-cluster approach requires more engineering to prevent points from incorrectly supporting an edge and disambiguation between fitting a line or a low-curvature edge. We therefore adopt the simpler cluster-then-fit strategy which is detailed next.
Clustering points into edges.  Graph traversal is a common strategy to cluster points where points form the vertices of the graph and an edge exists between two vertices only if they are spatially close [51, 18]. Clustering the points into their supporting edges then amounts to solving a graph traversal problem. This simple solution is usually made robust by integrating a smoothness constraint between neighboring points [53]: two vertices are connected not only if they are spatially close but if their direction is also similar. This strategy is used in [41] where the point’s direction is derived after training the edge field with a subsequent optimization whereas our method is simpler in that it learns both the edge point position and orientation directly.
Given the trained Gaussians, we define the graph as follows. The vertices are the oriented edge points: the vertex’s position is the Gaussian’s center and the vertex’s direction is the Gaussian’s principal direction. Vertices are neighbors if they are spatial neighbors and if they have similar orientations. A cluster is initialized with a point and a new point is added to the cluster if it is a neighbor of the last added point and if its orientation aligns with the orientation of the edge grown so far. To avoid adding points that do not lie on the edge but are close and oriented parallel to the edge, we also check if the new point’s orientation aligns with the line joining the last added point and the new one. For both orientation tests, we use a single orientation threshold $\theta$. The graph traversal results in clusters of points belonging to the same edge. We then fit edges to the resulting clusters in the form of line segments and cubic Bézier curves.
@@
Parametric Edge Fitting.  Once the points are clustered into individual edges, we next select whether a line or a cubic Bézier curve best fits the set of points. One solution [41] consists is first fitting lines until no more lines can be fitted with a geometric error below a given threshold then fitting curves to the remaining points. However, with edges of different lengths and different noise levels, tuning the geometric error threshold is complex and can result in many curves incorrectly modeled as lines, as observed in the EMAP [41] results. Instead, we propose to compare both the line and the curve model for each point cluster and select the best based on their relative geometric errors.
Let $e_{\text{c}}$ and $e_{\textit{l}}$ be the average residual error of the curve and line models respectively. We select the curve model if the error $e_{\text{c}}$ is smaller than a fraction $\delta$ of the line residual error, i.e., $e_{\text{c}}\leq\delta e_{\textit{l}}$. The parameter $\delta$ can be tuned to control the fraction of lines and curves in the output. As shown in the qualitative results (Fig. 4), this strategy is more effective than the ones adopted in NEF [83] and EMAP [41]. NEF [83] first approximates all edges with 2-control-points Bezier curves, i.e., 3D lines, and uses the estimated control point to further optimize the models into 3D curves when needed. This approach produces several degenerate curve configurations, especially close to the corners.
@@ 4.1 Experimental Setup
Datasets.  The quantitative evaluation is performed on the subsets of ABC-NEF [83] and the DTU [31] datasets used by EMAP [41] for evaluation. For fairness, we train on the 2D edge maps  [66, 57] they provide and use their released evaluation code. Qualitative results on selected scenes from Replica [65] and TnT [36] datasets are provided in the supplementary material.
@@
The ABC-NEF [83] subset of ABC dataset [38] is curated to evaluate 3D edge mapping. It consists of CAD models, ground truth parametric edges and 50 views rendered from around the object. In EMAP [41], models with cylinders and spheres are removed as they are inconsistent shapes for the task of edge mapping. Similarly, we use the 6 scenes from the DTU [31] dataset that contains multi-view images of everyday objects captured from fixed frontal viewpoints. A pseudo-ground truth of “edge-points” is created by projecting the dense 3D points reconstructed by a structured-light scanner [1] onto the 2D edge maps. Although we perform at par or better than the baselines on DTU, we observe that the pseudo-ground-truth introduces biases in the evaluation and we illustrate them in the qualitative results. Finally, qualitative results on selected scenes from Replica [65] and TnT [36] datasets are provided in the supplementary material.
@@
Figure 5: Qualitative results on DTU [31]. Comparison of the proposed 3D reconstruction method against baselines. NEAT [79] produces partially complete reconstructions with lines only. The edges from EMAP [41] are more complete but there are duplicate edge predictions for a single target edge in 3D, while our reconstruction is much cleaner with mostly a single predicted edge boundary (see zoom-ins).
Baselines.  We evaluate the line-SfM-based method LIMAP [43] that sets the state-of-the-art in geometry-based 3D line reconstruction and the 3D wireframe learning method NEAT [79]. We also evaluate the state-of-the-art 3D line / curve mapping methods NEF [83] and EMAP [41]. Following [41], NEAT [79] is not evaluated on ABC-NEF [83]: the authors observe that it often fails to train on the textureless renderings of the CAD models. The qualitative results we report are generated with the code and the weights released by the authors of the various baselines. For EMAP [41], at the time of writing, running the released code with the recommended configurations did not yield the same results as the ones shown in the paper or the website. We therefore report the results run on checkpoint versions graciously provided by the authors. The quantitative results for the baselines are reported as is from [41].
Metrics.  For quantitative evaluation, the process defined in [41] samples points along the predicted parametric edges and compares those points against points sampled at the same resolution on the ground-truth edges to compute the metrics. The accuracy defines the mean distance from the predicted points to the closest ground-truth points, and the completeness defines the mean distance between the ground-truth points and their nearest predicted point. For these two metrics, the lower the better. The precision at a distance threshold $\tau$ (P($\tau$)) measures the percentage of predicted points with at least one ground-truth point within distance $\tau$. Symmetrically, the recall (R($\tau$)) measures the percentage of ground-truth points for which a predicted point is within the distance threshold $\tau$. For precision and recall, the higher the better. We report the metrics for $\tau$ in 5, 10, and 20 millimeters (mm). We also report the training times of the neural implicit representations to show the advantage of the explicit representation adopted in this paper.
Although these metrics provide a meaningful evaluation of the reconstruction quality, they do not account for all the performance aspects of the 3D edge reconstruction. For example, duplicate edge reconstructions that lie close to a ground-truth edge are not penalized. This is observed in several reconstructions from LIMAP [43] and EMAP [41]. Further, this is especially important on the DTU dataset where pseudo-ground-truth 3D edge points do not lie exactly on the 3D edges but close to them (See Fig. 6).
Implementation details are provided in the suppl. mat.
4.2 Results
Figure 6: Qualitative results and pseudo-ground-truth on DTU [31]. We visualize our method (red) and the pseudo-ground-truth edge points (green) generated in [41] by projecting ground-truth 3D points onto edge images and using those projecting on edges in a majority of views [10]. Note that our method gets penalized in precision when it predicts geometrically correct 3D edges that are missing from the pseudo-ground-truth, e.g., the top of the smurf and the howl. Also, the pseudo-ground-truth contains thick patches of points that make it difficult for well-predicted edges to get a high recall. Still, our method faithfully reconstructs the 3D parametric edges of the objects.
@@ 4.2 Results
Our method produces the most complete results while being almost as precise as the methods leading that field. LIMAP [43] demonstrates impressive precision while missing curves and therefore lacking completeness. EMAP [41] performs better than our method under the 5mm error threshold although this trend is inverted under the 10mm and 20mm error thresholds. The performance difference under the 5mm threshold can be attributed to a combination of two things. Firstly, EMAP performs an extra point refinement step, that our method does not. Secondly, there is a bias introduced by the thickness of 2D edge maps generated by the detectors [66, 57]: we visually verify it by observing that the ground truth edges project close to the thick edges produced by these detectors, instead of projecting to its center. Shifting or scaling the Gaussians appropriately to counter this bias is left for future work.
@@
Evaluation on DTU [31] The quantitative results are reported in Tab. 2-left and the comparative qualitative results are shown in Fig. 5. While our method either outperforms or is at par with the baselines as per the defined quantitative evaluation, we observe biases in the evaluation due to the pseudo-ground-truth generation process. The pseudo-GT obtained by filtering out 3D points that do not project on 2D edge maps in enough views has two flaws. Firstly, the pseudo-GT points do not span the part of the scene that is not covered by a sufficient number of cameras. This issue is highlighted in  Fig. 6 where the pseudo-GT points in green show an incomplete coverage of the structure while our method correctly predicts a set of edges that completely cover the scene. Secondly, the 3D points that get labeled as pseudo-ground-truth edge points are not only the points lying on the ground-truth edges but also the points within an $\epsilon$-bound on the edge, and the range of the bound is a function of the 2D edge map thickness and the distance between the object and the camera. The first bias decreases the precision score for a set of edges that actually completely cover the scene as illustrated by our precision scores in Tab. 2. As for the second bias, i.e., the bias of thick edges, it not only penalizes the recall but it also promotes the presence of duplicate edges, as estimated by EMAP [83]. This can be seen in the examples shown in Fig. 5. Regarding the runtime, our method is as efficient as on the ABC-NEF [83] dataset and runs 22 times faster than NEF [41] and more than a hundred times faster than EMAP [41] (Tab. 2-right).
5 Limitations and Future work
The method inherits weaknesses of the original 3DGS [34] approach - it requires tuning of several parameters for the adaptive density control. Also, while the geometric regularization and clustering work robustly for simple CAD objects, they sometimes lead to artifacts and inaccuracies in more complex scenes. A few such issues are presented in the supplementary material. Our clustering method is driven by heuristics and therefore occasionally fails in certain parts of the scene. Incorporating structural priors can make this component more robust.
@@ 6 Conclusion
In this work, we propose a method for 3D edge reconstruction from images that explicitly learn the 3D edge points on top of which the 3D edges are fitted. The explicit representation is simple and casts 3D edge points as 3D Gaussians and the edge direction as the principal axis of the Gaussians. Such a representation allows for efficient rendering-based training supervised with off-the-shelf 2D edge maps. Results show that the proposed method is several times faster than the previous learning based approaches, while being slightly better or at-par in terms of the accuracy and completeness of estimated 3D edges. While the off-the-shelf 2D edge maps make for a relevant supervisory signal, we observe that they can introduce bias in the training or the evaluation, which calls for investigating better supervision in the future.
Acknowledgements.  This work was supported by the Czech Science Foundation (GACR) EXPRO (grant no. 23-07973X), the Chalmers AI Research Center (CHAIR), WASP and SSF. The compute and storage were partially supported by NAISS projects NAISS-2024/22-637 and NAISS-2024/22-237 respectively.
[…]
@@
Appendix A provides implementation details about the training of the edge-specialized Gaussian Splatting. Appendix B shows qualitative results over the scenes from the Replica [65] dataset used by the authors of EMAP [41] and three scenes from the Tanks and Temples dataset [36]. Appendix C discusses some limitations and failure cases of our method, pointing to relevant future work.
Appendix A Implementation details
Initialization.  Gaussian Position: For scenes from the DTU [31], Replica [65] and Tanks and Temples [36] datasets, we use the SfM [63] points as initialization. Note that random point initialization also produces reasonable, but slightly worse results. For ABC-NEF [83], we initialize our method with Gaussians centered at 10000 points randomly sampled in a unit cube. This is because the dataset comprises texture-less objects for which SfM [63] generates extremely sparse or no point reconstruction at all. Gaussian Scale: We use a constant initial value of 0.004 for all datasets. However, a point-dependent value based on the complexity of the neighboring region may be more robust. Gaussian Opacity: We use a constant initial value of 0.08 for all datasets. Gaussian Orientation: Random unit quaternions are used as initial values for all Gaussians.
Training.  We train the model for 500 epochs. For the first 30 epochs, we only train the position parameters so that the scale and the orientation of the Gaussian do not compensate for its incorrect position during rendering. Thus the training constrains the Gaussian’s position, i.e., its mean, to lie on 3D edges. We cull the Gaussians based on opacity and duplicate the ones with high positional gradients at regular intervals as in the original work [34]. The learning rates of different parameters are as follows. Position: starting with $1e^{-3}$, scaled with a factor of 0.75 every 10 epochs, 5 times. Scale: $2e^{-4}$ constant. Opacity: $3e^{-2}$ constant. Orientation: $1e^{-3}$ constant.
We use $k=4$ as the number of nearest neighbors for computing $\mathcal{L}_{\text{orient}}$ defined in Eq.(4) of the main paper. The weights of the loss function in Eq.(6) of the main paper are $\lambda_{1}=0.1$ and $\lambda_{2}=0.1$ for object level scenes from ABC-NEF [83] and DTU [31], while for larger scenes we use smaller values of $\lambda_{1}=0.01$ and $\lambda_{2}=0.01$. The geometric regularization assumes that the Gaussians are already positioned close to the edges, therefore we start applying this regularization at epoch 300. Note that the computation of nearest neighbors, required for the geometric regularization is computationally intensive and we observe that it is sufficient to only apply this regularization once in every 10 steps of the training process.
Clustering  During clustering, the alignment threshold is $\theta=0.8$ on ABC-NEF [83], which have clean straight lines, and $\theta=0.6$ on DTU [31] to account for the higher curvature of the shapes. During the parametric edge fitting, we fit a curve whenever the curve residual error is $\delta=0.5$ lower than the line residual error. For objects from the DTU dataset [31], due to the prior knowledge that the objects have more curves than lines, a larger $\delta=1$ is used.
For any further clarifications, please refer to the code released at https://github.com/kunalchelani/EdgeGaussians.
Appendix B Additional Qualitative Results